OpenStudy (anonymous):

Which value is a soluation of the equation 3-2x=-5 A.2 B.4 C.-2 D.1/5

OpenStudy (shamil98):

3-2x = -5 -3 -3 -2x = -8 divide by -2 both sides x = 4